# Universitat Autònoma de Barcelona, Catalunya

### v1

@@ Abstract
Transfer Learning Optical Character Recognition
@@ 1 Introduction
Acquiring large annotated corpora for certain languages is significantly more challenging than for plain Latin alphabets and, in some cases, outright prohibitive. This challenge becomes even more pronounced in the context of historical document analysis, where the scarcity, sparsity, and unique characteristics of the data render the annotation process both costly and, at times, unfeasible.
@@
As shown in Figure 1, this work leverages model merging algorithms to develop reading systems that can efficiently learn new alphabets—such as those in historical ciphered texts or low-resource languages—by learning robust features. Our approach enables data-efficient incorporation of non-Latin scripts without requiring centralized data (i.e., domains can be trained independently). Unlike [32], we adopt a meta-learning-inspired strategy to handle target datasets $g$ that differ significantly from the original distribution. We show that this improves adaptability in scenarios involving disjoint data distributions $d_{n}$.
@@
    A unified evaluation baseline is established across 20 major OCR datasets—including handwritten, scene, printed, historical, ciphered, and cross-lingual texts. Over 100 trained and fine-tuned models (baselines, per-language/cipher ablations, out-of-domain merges) are provided to the Document Analysis community. Their performance is documented in a publicly available model card [^1] https://eaudedata.github.io/ODAOCR/model_card/model_card.html, where all model weights can be accessed.
@@ 2 Related Work
As noted in [46], we can find examples of few-shot tasks as a semi-supervised pseudo-labeling task. These approaches range from pseudo-labeling for data augmentation [47, 55] to some curriculum-based approaches [8, 24]. Authors at [46], utilize a similar prototyping technique to alleviate the scarcity of handwritten document data. In this work, we aim to propose an optimization technique to improve recognition in low-resource scenarios, as the ones to be found in [46], without any additional mechanism than gradient descent algorithms. More specifically, we propose adapting narrow domains (historical ciphers and other low-resource alphabets) by generalizing on a wider set of tasks.
@@ 3.2 Optimization
This work demonstrates the capabilities of fused models in operating out-of-domain. Following the notation used in [15] and Section 3.1, we consider a model as vector of parameters $\theta\in\mathbb{R}^{m}$. Given a pretrained model, $\theta_{0}$, we fine-tune it with a training set $D=\{d_{1},d_{2},\ldots,d_{n}\}$, which converges towards $\theta^{D}_{T}$ (see Figure 2(a)), analogously, a sub-domain $d_{n}\in D$ optimizes models that converge towards $\theta^{d_{n}}_{T}$.
@@ 4 Experimental Setup
(k) Korean-[31]
(l) Bangla-[31]
(m) Arabic-[31]
(n) Chinese-[31]
(o) Hindi-[31]
(p) Japanese-[31]
Figure 3: Landscape of some of the datasets used in this research: Esposalles (a), George Washington (b), IAM (c), Parzival (d), Historical Maps (e), Borg Cipher (f), Vatican Cipher (g), Copiale Cipher (h), WordArt (i), AMR (j), MLT (k to p). Additionally, we present experiments on latin examples of HierText[26], SROIE[19], TotalText [9], and TextOCR[42].
@@ 4.1 Datasets
For in-domain training, we primarily use the domain of handwritten (HW) and scene text (Scene) recognition, on which we list Esposalles [37], IAM [27], George Washington [11] and Parzival (Figures 3(a) to 3(d)) for handwritten and CoCoText [50] and TextOCR [42] for scenes. Additionally, MLT19 [31] is considered as training set of the scene domain in its Latin subsampling; while the rest of splits (Arabic, Chinese, Japanese, Korean, Hindi and Bangla) in MLT19 are kept as fine-tuning data for transfer learning experiments (see Figures 3(m), 3(n), 3(p), 3(k), 3(o), 3(l)). SVT [53] and IIIT5K [29] are kept as evaluation data for scene text recognition as common practice in the literature [7].
Other data considered to experiment with the incorporation of new scripts are ciphered historical documents: Vatican [13], Copiale [21] and Borg [2] (see Figures 3(g), 3(h), 3(f)); which constitute the biggest challenge since it contains unseen handwritten symbols (with the exception of Vatican, containing a mixture of numbers, symbols and special accents) without further contextual information, such as background or incidental correlations, to rely on.
Importantly, we define a set which we consider “out-of-domain” for different reasons. Starting with printed documents, historical maps [54] are used as samples containing information not strictly related with text (see Figure 3(e)). Other kind of printed documents, such as receipts in SROIE [19] or forms in FUNSD [17] face challenges associated with arbitrary segments of text (such as named entities, codes or numbers), which, as noted by [52], reduces the reliance on shortcuts with respect data in $D$, on which one can associate words to real-world symbols or to build an implicit language model, which is the case for [37] with sentences having a fixed structure. Similarly, the AMR dataset [22] contains digits from domestic gas company meters, which contains arbitrary sequences combined with noise and flares (see Figure 3(j)). Lastly, TotalText[9] and WordArt datasets (see Figure 3), contain curved texts with artistic fonts. We consider this an important domain shift because ViTs are trained upon the assumption of vertically aligned text; on which the positional encoding plays a vital role.
@@ 4.2 Architecture Details
To ensure proper comparisons, we establish a common baseline across various in-domain training sets (handwritten and scene) and evaluation splits (out-of-distribution). We utilize a ViT encoder as described in [3]. However, we enhance the performance of the model by incorporating an additional encoding layer on top of the previous output, as suggested in [10]. This additional layer includes a positional encoding and is conditioned to a projection of the non-positionally-encoded input.
@@
Building upon previous literature [5, 10, 39], we train the transformer using a CTC [12] loss function. Unless stated otherwise, including pretraining, it is assumed to use 75 epochs of optimizing with Adam, a learning rate of $10^{-5}$ and 128 batch size. In the context of transfer learning to new languages, 30 epochs are used instead in order to promote models that can easily incorporate new knowledge; this hyperparametrization is empirically defined to maintain a common regime to all models while ensuring convergence in all cases.
@@ 5 Results and Discussion
Figure 5: Accuracy gain with respect baseline ($y=1$) in new languages (5(a)) or ciphered texts (5(b)).
@@
Avg (HW, Scene): In this approach, two models are independently trained on handwritten and scene datasets ($d^{hw},d^{scene}$). The resulting task vectors are then applied to the zero-shot model via averaging.
Avg (Ind.): In this case, task vectors are obtaining after dis-jointly training a different model for each training dataset ($\{d_{1},\ldots,d_{7}\}$). We then apply those vectors by averaging them to obtain the final model, which results in $\theta^{ind}$.
@@
Regarding the incorporation of new languages, after performing model merging, we analyze the impact of using $\theta^{ind}$ as an initialization point for learning previously unseen scripts in a new training stage. As shown in Figure 5(a), fine-tuning from the averaged models significantly accelerates the incorporation of new languages, in some cases achieving nearly three times the improvement compared to the baseline. In contrast, models trained directly on in-domain data within a centralized framework rarely outperform the baseline. Even when slight improvements occur, our distributed strategy consistently yields superior results, even when utilizing the same training data.
Moreover, in Figures 5(a) and 5(b), we valiate our design choice of using $U(Z,\theta)$ as seed for our experiments. The reason for the necessity of using a reasonably good seed for performing generalization, rather than a random seed could be attributed to the notion of locality observed in [15]. This is, randomly initialized models tend to significantly diverge when training on distributed data. Even though our strategy dominates distributed learning from a random seed; the last, tends also to outperform centralized approaches, which proves to be the least practical approach.
However, this scenario assumes that $\theta^{ind}_{1}$ is a converged model across all datasets after performing task arithmetic. This raises questions regarding both the data used and the decision to use task arithmetic ($T=1$) instead of a more extended meta-learning procedure. In what follows, we address these issues by taking a closer look at the behavior of such generalizing models.
@@
As it was previously stated, in this section we aim to address two questions regarding the effectiveness of this method for incorporating new scripts in character recognition systems. First, whether the learnt features of generalist models is independent from the extensiveness of the meta-learnt process or rather the models perform independently of its computing expense. As shown in Figure 6, we examine the average accuracy and standard deviation of fine-tuned models for both non-western scripts and historical ciphers across different sub-sampling percentages. We observe that performance converges upon reaching a critical data threshold (approximately 60% of $d_{n}$ for each sub-domain). However, at any sub-sampling percentage, our method consistently outperforms centralized pre-training. For a comprehensive version of this experiment, refer to the previously mentioned model card, where the results for each individual target dataset $g$ are detailed. From this analysis, we observe that the experiments in this section provide evidence that task arithmetic becomes more prominent at large pre-training scales. However, it is important to note that even in smaller sub-sampling regimes, the method consistently outperforms a centralized pre-training approach. This suggests that the benefits of task arithmetic extend beyond just large-scale scenarios, offering advantages even when data availability is limited.
Secondly, we inspect up to which degree the number of meta-learning steps $T$ can fatherly improve the results with respect using task arithmetic ($T=1$). In other words, it is explored if there is a limit for generalization on meta-learning regimes. Figure 7 shows the achieved accuracy after pre-training on $T=\{1,3,5,15,25,75\}$ and $k=\{75,25,15,5,3,1\}$. It is ensured to maintain a ratio of $k\times T=75$ to guarantee fair comparisons. From the computational point of view, every trained model performed the same amount of backward / forward passes on $D$, but distributed under a different amount of meta-learning steps ($T$). From the analysis conducted in Figure 7 we conclude that task arithmetic is not all you need, but rather a close version of the ideal case of meta-learning during a few distributed rounds. After performing many meta-learning steps (i.e. few steps on $U^{k}$ since $T=75$ and $k=1$) the performance decays due to an over-specialization of the model on the original source domains.
@@
Computational Resources
For practitioners aiming to fine-tune a pre-trained model, the estimated computational requirements are approximately 1 week of pre-training on a single RTX 3090 GPU, followed by around 5 days for training on sub-domains in a single-threaded, sequential manner. Model averaging takes roughly 1 minute on a CPU, while transfer learning requires about 5 hours. This total time can be significantly reduced by leveraging the distributed nature of our method, particularly by utilizing multiple GPUs in parallel during the sub-domain specialization stage. Additionally, all models are accessible in the provided download page.
@@ 6 Conclusions
In summary, as a take-home message for OCR generalization, we propose that practitioners begin by creating a subsample of the pre-training dataset, which is agnostic of specific modalities or organizational boundaries, but shares a common high-resource alphabet. This approach enables the training of multiple expert models on distributed subsets, which collectively encompass the same diversity as the original dataset $D$. The aggregated expertise of these models is particularly effective in out-of-domain scenarios, allowing the averaged model to better handle unannotated documents with significant distribution shifts. Furthermore, if these documents contain alphabets that differ from those in the training data, a brief period of fine-tuning with minimal annotations and a few epochs can rapidly adapt the model to the new conditions

### v2

# The OCR Quest for Generalization: Learning to recognize low-resource alphabets with model editing

@@ Abstract
Transfer Learning , Optical Character Recognition , Low-Resource Document Analysis
journal: Pattern Recognition
[…]
organization=Centre de Visió per Computador (CVC),
[…]
organization=Computer Science Department, Universitat Autònoma de Barcelona,
@@ 1 Introduction
Acquiring large annotated corpora for certain languages is significantly more challenging than for plain Latin alphabets and, in some cases, outright prohibitive. This challenge becomes even more pronounced in the context of historical document analysis, where scarce and unique data make annotation especially costly.
@@
As shown in Figure 1, this work leverages model merging algorithms to develop reading systems that can efficiently learn new alphabets—such as those in historical ciphered texts or low-resource languages—by learning robust features. Our approach enables data-efficient incorporation of non-Latin scripts without requiring centralized data (i.e., domains can be trained independently).
Unlike [37], we adopt a meta-learning-inspired strategy to handle target datasets $g$ that differ significantly from the original distribution. We show that this improves adaptability in scenarios involving disjoint data distributions $d_{n}$.
@@
    A unified evaluation baseline is established across 20 major OCR datasets - including handwritten, scene, printed, historical, ciphered, and cross-lingual texts. Over 100 trained and fine-tuned models (baselines, per-language/cipher ablations, out-of-domain merges) are provided to the Document Analysis community. Their performance is documented in a publicly available model card [^1] https://eaudedata.github.io/ODAOCR/model_card/model_card.html, where all model weights can be accessed.
@@ 2 Related Work
As noted in [51], we can find examples of few-shot tasks as a semi-supervised pseudo-labeling task. These approaches range from pseudo-labeling for data augmentation [52, 61] to some curriculum approaches [8, 28]. Authors at [51], utilize a similar prototyping technique to alleviate the scarcity of handwritten document data. In this work, we aim to propose an optimization technique to improve recognition in low-resource scenarios, as the ones to be found in [51], without any additional mechanism than gradient descent algorithms. More specifically, we propose adapting narrow domains (historical ciphers and other low-resource alphabets) by generalizing on a wider set of tasks.
@@ 3.1 Notation
3.2 Preliminaries
In this section, we aim to establish the necessary preliminaries to unify the notation used throughout this work, given the diverse research directions that converge into the FedAvg algorithm [32], it is necessary to establish equivalences between Reptile [37], task arithmetic [16] and FedAvg, as all of the stated methods induce an interpretation of a common algorithm from different perspectives.
Intuitively, this methods train different models on different data distributions (see blue dots in Figure 1). While the data separation usually consists on disjoint distributions, in the case of Reptile [37] the data distributions are randomly sampled from an original dataset where one category is left out for evaluation; which stabilizes the training process by merging models that did not diverge too much. This models are then arithmetically merged [16] for conducting posterior training stages. In our case, we go one step further and utilize this FedAvg models to conduct fine-tunning in a domain adaptation set-up (see yellow dot in Figure 1).
First, the FedAvg algorithm [32] initializes a network with a set of parameters $\theta$ and data $D$ that is disjointly distributed among $N$ clients $\{d_{1},d_{2},\ldots,d_{n}\}$. In real-world applications, each client represents a different dataset of a common task. Once the different partitions are defined, an update function $U(D,\theta)$ is introduced, which consists of $k$ stochastic optimization steps (such as Adam [23]) using the data partition $d_{n}$. This update process results in $U^{k}(d_{n};\theta)$. A learning optimization process, defined in Equation 1, is then repeated $T$ times. This training regime is ruled by the number of stochastic optimization steps $k$ and the number of aggregation steps $T$.
$\theta_{t+1}\leftarrow\theta_{t}+\frac{1}{N}\sum^{N}_{n}\left[U^{k}(d_{n};\theta_{t})-\theta_{t}\right]$ (1)
As noted in [11] and [19], this optimization mechanism is equivalent to Reptile [37], which samples $\{d_{1},d_{2},\ldots,d_{n-1}\}$ partitions disjoint to $d_{n}$ from Omniglot [25] or Mini-ImageNet [56], leaving a category out for meta-learning evaluation. Authors then try to few-shot the categories found in the later partition, which are originally contained in $D$.
Note that when $T=1$ and, for any partition, $k>\left|d_{n}\right|$, the averaging of model differences is performed just once, but multiple updates on the dataset $d_{n}$ are conducted. This process creates an ensemble of specialist models that are subsequently merged into a generalist one. In this scenario, authors in [16] consider $\tau=U^{k}(d_{n};\theta_{0})-\theta_{0}$ a task vector. Additionally, in [40], it is highlighted how an overspecialization ($k\gg\left|d_{n}\right|$) can lead to bad generalization due to lose of “linearity" of the training regime.
In this work, we initialize $\theta_{0}$ by training on a Latin dataset $Z$. This ensures that $\theta_{0}=U(Z,\theta)$ does not diverge significantly when updating on different domains in Latin examples for handwritten and scene text recognition, which we consider it to be our source dataset $D$. In this way, as recommended by [40], we ensure that locality is maintained during the meta-learning process. This is something that Reptile addresses by stating an scenario where the different distributions are sampled by an original one, therefore models do not require to accumulate very different visual features.
Because our scenario should be accountant of the inherent constraints of Document Intelligence (security, sparsity, unbalance…), we advocate for the Federated set-up. This ensures a transferability inherent to Reptile while maintaining the datasets disjoint and, therefore, ensuring privacy. For doing so, it is necessary to start from a reasonable seed in a task-arithmetic manner.
In contrast to [37], we utilize meta-learning to adapt to data that strongly differs from the original distribution, this is indicated by a target dataset $g$. While current research focused on evaluating model-merging either in zero-shot or few-shot in-domain examples; we demonstrate how federated learning can enhance reading systems that need to adapt to scenarios such as historical ciphered alphabets or low-resource languages while retaining distributed advantages in a setting where the different partitions $d_{n}$ originate from disjoint data distributions. In the following sections, we will describe the characteristics of these partitions.
@@ 3.3 Optimization
This work demonstrates the capabilities of fused models in operating in transfer-learning scenarios. Following the notation used in [16] and Section 3.1, we consider a model as vector of parameters $\theta\in\mathbb{R}^{m}$. Given a pretrained model, $\theta_{0}$, we fine-tune it with a training set $D=\{d_{1},d_{2},\ldots,d_{n}\}$, which converges towards $\theta^{D}_{T}$ (see Figure 2(a)), analogously, a sub-domain $d_{n}\in D$ optimizes models that converge towards $\theta^{d_{n}}_{T}$.
@@ 4 Experimental Setup
(k) [21]
(l) Korean-[36]
(m) Bangla-[36]
(n) Latin-[36]
(o) Arabic-[36]
(p) Chinese-[36]
(q) Hindi-[36]
(r) Japanese-[36]
(s) [9]
(t) [47]
Figure 3: Landscape of some of the datasets used in this research: Esposalles (a), George Washington (b), IAM (c), Parzival (d), Historical Maps (e), WordArt (f), AMR (g), Borg Cipher (h), Copiale Cipher (i), Vatican Cipher (j), SROIE (k), MLT (l to r), TotalText (s), and TextOCR (t).
@@ 4.1 Datasets
For in-domain training, we primarily use the domain of handwritten (HW) and scene text (Scene) recognition, on which we list Esposalles [42], IAM [31], George Washington [12] and Parzival (Figures 3(a) to 3(d)) for handwritten and CoCoText [55] and TextOCR [47] for scenes. Additionally, MLT19 [36] is considered as training set of the scene domain in its Latin subsampling; while the rest of splits (Arabic, Chinese, Japanese, Korean, Hindi and Bangla) in MLT19 are kept as fine-tuning data for transfer learning experiments (see Figures 3(o), 3(p), 3(r), 3(l), 3(q), 3(m)). SVT [59] and IIIT5K [34] are kept as evaluation data for scene text recognition as common practice in the literature [7].
Other data considered to experiment with the incorporation of new scripts are ciphered historical documents: Vatican [14], Copiale [24] and Borg [2] (see Figures 3(j), 3(i), 3(h)); which constitute the biggest challenge since it contains unseen handwritten symbols (with the exception of Vatican, containing a mixture of numbers, symbols and special accents) without further contextual information, such as background or incidental correlations, to rely on.
Importantly, we define a set which we consider “out-of-domain” for different reasons. Starting with printed documents, historical maps [60] are used as samples containing information not strictly related with text (see Figure 3(e)). Other kind of printed documents, such as receipts in SROIE [21] or forms in FUNSD [18] face challenges associated with arbitrary segments of text (such as named entities, codes or numbers), which, as noted by [58], reduces the reliance on shortcuts with respect data in $D$, on which one can associate words to real-world symbols or to build an implicit language model, which is the case for [42] with sentences having a fixed structure. Similarly, the AMR dataset [26] contains digits from domestic gas company meters, which contains arbitrary sequences combined with noise and flares (see Figure 3(g)). Lastly, TotalText[9] and WordArt datasets (see Figure 3), contain curved texts with artistic fonts. We consider this an important domain shift because ViTs are trained upon the assumption of vertically aligned text; on which the positional encoding plays a vital role.
@@ 4.2 Architecture Details
To ensure proper comparisons, we establish a common baseline across various in-domain training, evaluation and transfer learning splits. We utilize a ViT encoder as described in [3]. However, we enhance the performance of the model by incorporating an additional encoding layer on top of the previous output, as suggested in [10]. This additional layer includes a positional encoding and is conditioned to a projection of the non-positionally-encoded input.
@@
Building upon previous literature [5, 10, 44], we train the transformer using a CTC [13] loss function. Unless stated otherwise, it is assumed to use 75 epochs of optimizing with Adam, a learning rate of $10^{-5}$ and 128 batch size. In the context of learning new alphabets, 30 epochs are used instead in order to promote models that can easily incorporate new knowledge; this hyperparametrization is empirically defined to maintain a common regime to all models while ensuring convergence in all cases.
@@ 5 Results and Discussion
Table 3: Accuracy results for transfer learning across multiple languages using different strategies. We compare centralized fine-tuning and task arithmetic (distributed), both starting from a shared pretrained model ($Z$). We also evaluate task arithmetic from a randomly initialized model. Despite using the same data splits and number of examples, the distributed strategy consistently outperforms centralized fine-tuning when transferring to low-resource alphabetic systems.
Vatican Borg Copiale Arabic Chinese Japanese Korean Bangla Hindi ×$\times\Delta$
Baseline U ( Z ) .549 .382 .825 .131 .020 .116 .282 .260 .470 -
Centralized (from Z ) .465 .000 .794 .175 .0103 .0951 .214 .115 .384 0.69
Distributed (from Z) .591 .505 .838 .422 .073 .197 .374 .462 .514 1.62
Distributed (random) .480 .272 .930 .410 .010 .114 .271 .266 .282 1.04
5.1 Quantitative Analysis
@@
Avg (HW, Scene): In this approach, two models are independently trained on handwritten and scene datasets ($d^{hw},d^{scene}$). The resulting task vectors are then applied to the zero-shot model via averaging. Avg (Ind.): In this case, task vectors are obtaining after dis-jointly training a different model for each training dataset ($\{d_{1},\ldots,d_{7}\}$). We then apply those vectors by averaging them to obtain the final model, which results in $\theta^{ind}$.
@@
Regarding the incorporation of new languages, after performing model merging, we analyze the impact of using $\theta^{ind}$ as an initialization point for learning previously unseen scripts in a new training stage. As shown in Table 3, fine-tuning from the averaged models significantly accelerates the incorporation of new languages, in some cases achieving nearly three times the improvement compared to the baseline. In contrast, models trained directly on in-domain data within a centralized framework rarely outperform the baseline. Even when slight improvements occur, our distributed strategy consistently yields superior results, even when utilizing the same training data.
Moreover, in Table 3, we valiate our design choice of using $U(Z,\theta)$ as seed for our experiments. The reason for the necessity of using a reasonably good seed for performing generalization, rather than a random seed could be attributed to the notion of locality observed in [16]. This is, randomly initialized models tend to significantly diverge when training on distributed data. Even though our strategy dominates distributed learning from a random seed; the last, tends also to outperform centralized approaches, which proves to be the least practical approach.
However, this scenario assumes that $\theta^{ind}_{1}$ is a converged model across all datasets after performing task arithmetic. This raises questions regarding both the data used and the decision to use task arithmetic ($T=1$) instead of a more extended meta-learning procedure.
(a) Arabic
(b) Bangla
(c) Chinese
(d) Korean
(e) Japanese
(f) Hindi
(g) Vatican
(h) Borg
(i) Copiale
Figure 6: Performance of the different models trained on multilingual and ciphered datasets under different sub-sampling proportions of the original data.
@@
As it was previously stated, in this section we aim to address two questions regarding the effectiveness of this method for incorporating new scripts in character recognition systems. First, whether the learnt features of generalist models is independent from the extensiveness of the meta-learnt process or rather the models perform independently of its computing expense. As shown in Figure 5, we examine the average accuracy and standard deviation of fine-tuned models for both non-western scripts and historical ciphers across different sub-sampling percentages. We observe that performance converges upon reaching a critical data threshold (approximately 60% of $d_{n}$ for each sub-domain). However, at any sub-sampling percentage, our method consistently outperforms centralized pre-training. In Figure 6 we observe how some alphabets (such as Chinese or Borg) are on a need for task arithmetic in order to converge fast enough to yield any result.
From this analysis, we observe that the experiments in this section provide evidence that task arithmetic becomes more prominent at large pre-training scales. However, it is important to note that even in smaller sub-sampling regimes, the method consistently outperforms a centralized pre-training approach. This suggests that the benefits of task arithmetic extend beyond just large-scale scenarios, offering advantages even when data availability is limited.
Secondly, we inspect up to which degree the number of meta-learning steps $T$ can fatherly improve the results with respect using task arithmetic ($T=1$). In other words, it is explored if there is a limit for generalization on meta-learning regimes. Figure 7 shows the achieved accuracy after pre-training on $T=\{1,3\dots 75\}$ and $k=\{75,25\dots 1\}$. It is ensured to maintain a ratio of $k\times T=75$ to guarantee fair comparisons. From the computational point of view, every trained model performed the same amount of backward / forward passes on $D$, but distributed under a different amount of meta-learning steps ($T$). From the analysis conducted in Figure 7 we conclude that task arithmetic is not all you need, but rather a close version of the ideal case of meta-learning during a few distributed rounds. After performing many meta-learning steps (i.e. few steps on $U^{k}$ since $T=75$ and $k=1$) the performance decays due to an over-specialization of the model on the original source domains.
@@
For practitioners aiming to fine-tune a pre-trained model, the estimated computational requirements are approximately 1 week of pre-training on a single RTX 3090 GPU, followed by around 5 days for training on sub-domains in a single-threaded, sequential manner. Model averaging takes roughly 1 minute on a CPU, while transfer learning requires about 5 hours. This total time can be significantly reduced by leveraging the distributed nature of our method, particularly by utilizing multiple GPUs in parallel during the sub-domain specialization stage.
@@ 6 Conclusions
In summary, as a take-home message for OCR generalization, we propose that practitioners begin by creating a subsample of the pre-training dataset, which is agnostic of specific modalities or organizational boundaries, but shares a common high-resource alphabet. This approach enables the training of multiple expert models on distributed subsets, which collectively encompass the same diversity as the original dataset $D$. The aggregated expertise of these models is particularly effective in out-of-domain scenarios, allowing the averaged model to better handle unannotated documents with significant distribution shifts. Furthermore, if these documents contain alphabets that differ from those in the training data, a brief period of fine-tuning with minimal annotations and a few epochs can rapidly adapt the model to the new conditions.
The paper shows that task arithmetic and distributed training enable robust OCR across low-resource alphabets and historical ciphered texts, outperforming classical fine-tuning while retaining high in-domain accuracy.